OpenStudy (anonymous):

x(x-6)(x-9)=v in standard form pleaseeeee

jhonyy9 (jhonyy9):

(x^2 -6x)(x-9)=x^3 -6x^2 -9x^2 +54x =x^3 -15x^2 +54x - this will be the standard form

OpenStudy (anonymous):

thanks

jhonyy9 (jhonyy9):

yw good luck bye